OpenStudy (anonymous):

ALGEBRA 1, PLEASE HELP!! WILL BE AWARDED! Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (acxbox22):

perpendicular slope is the opposite reciprocal of the lines slope so the slope for the perpendicular line would be 1 because -1's opposite reciprocal would be 1

OpenStudy (acxbox22):

now we have a point and slope so we can put it in point slope then convert it to slope intercept

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but i don't understand that.

OpenStudy (acxbox22):

what dont you understand?

OpenStudy (anonymous):

how to solve the problem.

OpenStudy (acxbox22):

in point slope form here is the equation of the perpendicular line y+2=1(x+4) in slope intercept it would be y=x+2 <<<< that is your answer

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

could you help me with another one?

OpenStudy (acxbox22):

sure but remember to post each question separately, thanks

OpenStudy (anonymous):

okay